"Salt water" refers to a mixture in which salt has been dissolved in water, with
no salt crystals remaining in the bottom of the container. When evaporation takes place, only
water leaves the solution, leaving all the salt in the remaining solution, or "brine."
Suppose 100 cc of salt water are set out in the sun. The original solution was
99 % water. After a period of time a sufficient amount of water has evaporated so
that the brine solution is only 98 % water. How much water must have
evaporated?

Answer by TimothyLamb(4379)

(0.01)100 = (0.02)(100 - x)
1 = 2 - (0.02)x
(0.02)x = 1
x = 1/(0.02)
---
answer:
x = 50 cc
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php